### v1

# Optimal structure learning and conditional independence testing

Abstract
We establish a fundamental connection between optimal structure learning and optimal conditional independence testing by showing that the minimax optimal rate for structure learning problems is determined by the minimax rate for conditional independence testing in these problems. This is accomplished by establishing a general reduction between these two problems in the case of poly-forests, and demonstrated by deriving optimal rates for several examples, including Bernoulli, Gaussian and nonparametric models. Furthermore, we show that the optimal algorithm in these settings is a suitable modification of the PC algorithm. This theoretical finding provides a unified framework for analyzing the statistical complexity of structure learning through the lens of minimax testing.
Contact: minggao@chicagobooth.edu, yuhaowang@u.nus.edu, bryon@chicagobooth.edu
1 Introduction
Graphical models are important tools in machine learning for representing complex dependency structures that arise in a wide range of applications in causality, artificial intelligence, and statistics (Spirtes et al., 2000; Pearl, 2010; Murphy, 2012; Zhang et al., 2013). A crucial preliminary step when using graphical models is structure learning, which seeks to recover the underlying graph from data. The accuracy of structure learning directly impacts the performance of downstream tasks. It is well-known that structure learning is closely related to conditional independence (CI) testing, which itself is a fundamental topic that has attracted significant attention in recent years. Owing to its foundational role, CI testing has been extensively studied both methodologically and theoretically, leading to a rich body of literature on its statistical properties (Zhang et al., 2011; Shah and Peters, 2020; Canonne et al., 2018; Neykov et al., 2021). In particular, structure learning methods often rely on CI testing as a subroutine for inferring graphical dependencies (Spirtes and Glymour, 1991; Maathuis et al., 2018).
The relationship between CI testing and structure learning has long been understood from the algorithmic perspective, which takes CI tests as black-box oracles and often assumes perfect CI information to guarantee graph recovery. However, the relationship between their finite-sample complexity and statistical hardness has yet to be formalized. Specifically, given that graphical models are intrinsically defined by the Markov property, a deeper understanding of how the statistical difficulty of CI testing governs the learning of graphical structures is both natural and fundamental.
In undirected graphical models (Markov random fields), the information-theoretic limits of structure learning—specifically, the minimum number of samples required for reliable graph recovery—have been widely investigated  (Wang et al., 2010; Drton and Maathuis, 2017; Misra et al., 2020). For directed acyclic graphs (DAGs), however, much of the optimality literature has focused on the Gaussian setting, where the analysis can leverage strong distributional properties. As a result, existing optimality results do not readily extend to discrete or nonparametric settings. In contrast, CI testing in its own right has been studied more broadly, with well-established minimax results across various distributional settings (Canonne et al., 2018; Neykov et al., 2021). This contrast suggests an opportunity to leverage statistical results in CI testing to better understand the sample complexity of structure learning for DAGs in general.
In this paper, we establish a fundamental relationship between the minimax optimality of structure learning in DAG models and the minimax optimality of CI testing, going beyond the Gaussian setup to general statistical models. We focus on poly-forests, a tractable yet rich subclass of DAGs that captures structured dependencies (Chow and Liu, 1968; Dasgupta, 1999), serving as a principled starting point toward fully general DAGs. Although optimality for learning general DAGs has been studied (Gao et al., 2022), specific distributional assumptions are needed in the analysis (e.g. linear SEM with equal error variances). We develop results under generic conditions that apply to both parametric and nonparametric distributions. Our work provides a framework connecting the statistical complexity of CI testing with that of structure learning, offering new insights into the understanding of sample-efficient DAG recovery in diverse modeling settings.
1.1 Contributions
Our contributions in this work are twofold:
  1. 1.
    We establish a connection (Theorem 3.1) between the statistical complexity of conditional independence testing and structure learning problems. We show that the optimal sample complexity of learning any poly-forest is
    $n\asymp\frac{\log d}{c^{\alpha}}\,,$
    when the optimal testing radius of the corresponding CI testing problem is $n^{-1/\alpha}$ for some $\alpha>0$ depending on the modeling setup, and $c$ is a parameter that captures the signal strength (cf. Section 2).
  2. 2.
    We apply this general result to derive the minimax optimal sample complexity for several practical examples, including Bernoulli, Gaussian, and nonparametric continuous distributions, and discuss their differences. We also show that the optimal sample complexity is achieved by an efficient algorithm based on the classical PC algorithm.
In addition, we conduct experiments to verify our theoretical findings.
1.2 Related work
Conditional independence testing
CI testing forms a crucial building block in machine learning and reasoning tasks. Given a triplet of random variables $(X,Y,Z)$, the problem asks whether $X$ is conditionally independent of $Y$ given $Z$. Numerous methods (e.g. Dawid, 1979; Fukumizu et al., 2007; Ramsey, 2014; Li and Fan, 2020) have been developed to address this fundamental problem. One prominent class of methods relies on measuring the distance between the conditional distribution $P(X,Y\,|\,Z)$ and the product of the marginals $P(X\,|\,Z)P(Y\,|\,Z)$. For Gaussian variables, the problem reduces to testing whether the partial correlation is zero (Fisher, 1915). For discrete variables, hypothesis tests based on chi-squared statistics are widely employed Ireland and Kullback (1968); Darroch et al. (1980). Recent advancements have also explored kernel-based methods Gretton et al. (2007); Zhang et al. (2011) and information-theoretic measures such as conditional mutual information (Runge, 2018; Berrett and Samworth, 2019) that can capture complex nonlinear dependencies in a nonparametric fashion.
From the theoretical standpoint, CI testing has been investigated from the perspective of minimax optimality. For discrete distributions, the problem is well-understood and the minimax optimal rates have been established (Chan et al., 2014; Diakonikolas and Kane, 2016; Canonne et al., 2018). In the nonparametric setting, Neykov et al. (2021) studied the fundamental limits of CI testing, deriving minimax optimal rates under smoothness assumptions. In addition, practical CI tests have been proposed (Kim et al., 2022, 2024). In particular, Jamshidi et al. (2024) devised a Von Mises estimator for mutual information as a valid CI test, and showed it achieves the parametric rate once the density is sufficiently smooth. These theoretical results provide valuable insights into the inherent difficulty of CI testing under different distributional setups.
Graphical model structure learning
Learning the structure of graphical models from observational data is a fundamental problem in machine learning. For undirected graphical models, the problem reduces to support recovery of the precision matrix (MEINSHAUSEN and BÜHLMANN, 2006; Friedman et al., 2008; Cai et al., 2011; Liu et al., 2009). For DAG learning, approaches can be broadly categorized into three main paradigms: score-based methods (e.g. greedy equivalence search) (Chickering, 2002; Nandy et al., 2018), constraint-based methods (e.g. PC algorithm) (Spirtes and Glymour, 1991; Friedman et al., 2013), and methods that leverage specific distributional assumptions (Shimizu et al., 2006; Hoyer et al., 2008; Peters et al., 2011, 2014; Peters and Bühlmann, 2014). Especially, constraint-based methods rely on valid CI tests and the faithfulness assumption to operate (Kalisch and Bühlman, 2007; Marx et al., 2021), thus development of reliable and efficient CI tests directly impacts the performance of these structure learning methods. For tree-structured models, the Chow-Liu algorithm (Chow and Liu, 1968; Chow and Wagner, 1973) provides an efficient method to find the optimal undirected tree structure by constructing a maximum weight spanning tree based on pairwise mutual information. Learning poly-trees/forests is generally more complex while more tractable than learning general DAGs REBANE (1987); Srebro (2003); Tan et al. (2010, 2011). Recent developments include (Gao and Aragam, 2021; Azadkia et al., 2021; Jakobsen et al., 2022). Furthermore, learning poly-trees/forests is relevant in various applications, including reasoning and density estimation (Kim and Pearl, 1983; Liu et al., 2011).
Sample complexity of structure learning
Understanding the sample complexity of learning graphical model structures is crucial for assessing the feasibility and reliability of structure learning algorithms. For undirected graphical models, optimal sample complexities have been derived for various classes of undirected graphs, such as sparse graphs with bounded degree Wang et al. (2010); Santhanam and Wainwright (2012); Bresler (2015); Vuffray et al. (2016); Abbe (2018); Misra et al. (2020). In contrast, determining the sample complexity for DAG learning is considerably more challenging due to the inherent asymmetry and the larger space of possible graph structures. Existing work often rely on specific assumptions about the underlying data distributions. Ghoshal and Honorio (2017) provided lower bounds for structure learning of general DAG models. Gao et al. (2022) established matching upper and lower bounds on the sample complexity as $\Theta[q\log(d/q)]$ for learning Gaussian DAGs under the restrictive equal variance assumption (Peters and Bühlmann, 2014; Loh and Bühlmann, 2014; Chen et al., 2019), where $q$ is the degree of the DAG, and this optimality result has been further extended to linear dynamic models (Veedu et al., 2024). Wang et al. (2024) concluded the optimal sample complexity as $\Theta(\log d/c^{2})$ for learning Gaussian poly-trees, where $c$ is the faithfulness parameter, and provided an efficient algorithm based on classic PC algorithm (Spirtes et al., 2000). For nonparametric setting, Jamshidi et al. (2024) applied the devised Von Mises CI test to the PC algorithm and obtain a dependence of $\mathcal{O}[(\frac{\Delta}{I_{\min}}\log d)^{2}]$, where $\Delta$ is the graph degree and $I_{\min}$ is a lower bound on the conditional mutual information, assuming the density function is sufficiently smooth and lower bounded (from zero). In contrast, our work focuses on establishing a general connection between minimax CI testing and structure learning, including method-agnostic minimax lower bounds, that apply broadly to general distributions with minimal assumptions.
2 Preliminaries
Given a directed acyclic graph (DAG) $G=(V,E)$, $\operatorname{pa}(k)=\{j:(j,k)\in E\}$ denotes the parents of node $k\in V$. The skeleton of $G$, $\operatorname{sk}(G)$, is the undirected graph formed by removing directions of all the edges in $G$. A triplet $(j,\ell,k)$ is called unshielded if both $j,k$ are adjacent to $\ell$ but not adjacent to each other, graphically $j-\ell-k$; and is called a $v$-structure if additionally $j,k$ are parents of $\ell$, i.e. $j\rightarrow\ell\leftarrow k$. A path is a sequence of distinct nodes $(h_{1},\ldots,h_{\ell})$ such that $(h_{j},h_{j+1})$ is in $\operatorname{sk}(G)$. A forest is an undirected graph where any two nodes are connected by at most one path. A poly-forest is a DAG whose skeleton is a forest. Denote the set of all poly-forests over $d$ nodes as $\mathcal{T}=\mathcal{T}_{d}$.
A distribution $p$ satisfies the Markov property with respect to a DAG $G$ with $d$ nodes if the following factorization holds
$p(X)=p(X_{1},\dots,X_{d})=\prod_{k=1}^{d}p(X_{k}\,|\,\operatorname{pa}(k))$
We consider the problem of structure learning, with a focus on poly-forests, where we assume the existence of a poly-forest $G\in\mathcal{T}$ such that $p$ is Markov to $G$, and we aim to recover $G$ given $n$ i.i.d. samples from $p$. In general, it is well-known that the DAG is not identifiable from observational data alone. Assuming faithfulness, $G$ is identified up to its Markov equivalence class (MEC), which is the set of DAGs that encode the same set of conditional independencies as $G$ and is represented by completed partially directed acyclic graph (CPDAG), denoted by $\overline{G}$. We refer the readers to Koller and Friedman (2009) for more preliminaries on graphical models.
2.1 Measuring dependence in general models
Since we assume the underlying DAG belongs to the class of poly-forests, the usual faithfulness assumption can be relaxed to tree-faithfulness (Wang et al., 2024) for successful recovery. To derive uniform, finite-sample bounds, we also need conditions on the minimum signal strength, as is standard in model selection and testing literature. We first define a general notion of dependence measure that will be used to quantify the signal strength.
Definition 1 (Dependence measure).
Let $(X,Y,Z)\sim p$ be a triplet of random variables subject to some distribution $p$. A (conditional) dependence measure $m(X;Y\,|\,Z)$ is a functional of $p$ such that (1) $m(X,Y\,|\,Z)\geq 0$; and (2) $m(X;Y\,|\,Z)=0$ if and only if $X\perp\!\!\!\perp Y\,|\,Z$ under $p$.
If $Z=\emptyset$, we simplify the notation by writing the marginal dependence as $m(X;Y)=m(X;Y\,|\,\emptyset)$. This general measure of dependence is used to simplify our main theorem statement; in our examples (Sections 4-5), we specify the dependence measure as the usual correlation coefficient for Gaussian distributions, total variation with the product of marginals for Bernoulli distributions, and total variation distance to the nearest independent instance for nonparametric continuous distributions. Using this dependence measure, we can now define strong tree-faithfulness in a generic form.
Definition 2 ($c$-strong tree-faithfulness).
A distribution $p$ is $c$-strong tree-faithful to a poly-forest $G$ with respect to the dependence measure $m$ if
  1. 1.
    For any two nodes connected $j-k$, we have $m(X_{k};X_{j}\,|\,X_{\ell})\geq c$ for $\ell\in V\cup\{\emptyset\}\setminus\{k,j\}$;
  2. 2.
    For any $v$-structure $k\rightarrow\ell\leftarrow j$, we have $m(X_{k};X_{j}\,|\,X_{\ell})\geq c$.
2.2 CI testing and structure learning
Since we aim to establish a statistical connection between CI testing and structure learning, we define each problem and the associated statistical quantities of interest as follows. In particular, given the close relationship between these two problems and the fact that structure learning often relies on CI testing as a subroutine, we first introduce the CI testing problem before defining the poly-forest learning problem. To ground the abstract formulation of CI testing, we begin with a concrete example under the nonparametric setting to highlight the key concepts and notations. See Section 5 and Appendix F for more details of this example.
Example 1 (Nonparametric models).
Let $\mathcal{P}$ be all continuous distributions supported on $[0,1]^{3}$ that admit Lipschitz continuous densities $p$. For each distribution $p\in\mathcal{P}$ over a triplet of variables $(X,Y,Z)$, we measure the dependence between $X$ and $Y$ given $Z$ by the total variation distance $\inf_{q\in\mathcal{Q}}\|p-q\|_{1}$, where $\mathcal{Q}$ is the set of all conditionally independent distributions in $\mathcal{P}$. This serves as a valid dependence measure $m$ for nonparametric distributions. The nonparametric CI testing problem asks for a test to distinguish the following two hypotheses:
$\mathcal{H}_{0}:$ $\quad p(X,Y,Z)\in\mathcal{P}\quad\text{ s.t. }\quad X\perp\!\!\!\perp Y\,|\,Z$
$\mathcal{H}_{1}:$ $\quad p(X,Y,Z)\in\mathcal{P}\quad\text{ s.t. }\quad\inf_{q\in\mathcal{Q}}\|p-q\|_{1}\geq c\,,$
for some $c>0$. A sufficiently large $c$ is necessary for consistent testing (Shah and Peters, 2020), and is also used to study the statistical hardness in terms of the minimax testing radius (defined in the sequel).
Building on the intuition from Example 1, we formally introduce the CI testing problem in a generic form, which can be instantiated under various distributional settings. Based on the elements of CI testing, we will then proceed to define the poly-forest learning problem such that the relationship between the two problems is explicit and clear. At its core, CI testing is a fundamental statistical problem concerned with distinguishing distributions over triplet of variables $(X,Y,Z)$.
Definition 3 (Conditional independence testing).
A conditional independence testing problem $\mathcal{C}(\mathcal{P},m,c)$ is defined by a model class of distributions $\mathcal{P}$ and dependence measure $m$, and aims to distinguish two hypotheses of distributions:
$\mathcal{H}_{0}:$ $\quad p(X,Y,Z)\quad\text{ s.t. }\quad m(X;Y\,|\,Z)=0\iff X\perp\!\!\!\perp Y\,|\,Z$
$\mathcal{H}_{1}:$ $\quad p(X,Y,Z)\quad\text{ s.t. }\quad m(X;Y\,|\,Z)\geq c$
where $p\in\mathcal{P}$. A conditional independence test $\psi$ is a function of data $\{X^{(i)},Y^{(i)},Z^{(i)}\}_{i=1}^{n}\mapsto\{0,1\}$; i.e. $\psi=0$ indicates $X\perp\!\!\!\perp Y\,|\,Z$. The minimax optimal testing radius of $\mathcal{C}(\mathcal{P},m,c)$ is the infimum of $c=c_{n}>0$ in terms of sample size $n$ such that the following holds:
$\inf_{\psi}\bigg{\{}\sup_{p\in\mathcal{H}_{0}}\mathbb{E}_{p}[\psi]+\sup_{p\in\mathcal{H}_{1}}\mathbb{E}_{p}[1-\psi]\bigg{\}}\leq\frac{1}{10}\,.$
The number $1/10$ here can be replaced with any small constant less than one. For a certain CI testing problem, one needs to specify the model class $\mathcal{P}$, and the dependence measure $m$. Consequently, the hypothesis classes $\mathcal{H}_{0},\mathcal{H}_{1}$ are determined. In Example 1, $\mathcal{P}$ includes all the Lipschitz densities over $[0,1]^{3}$ and $m$ is given by the total variational distance. One goal of studying $\mathcal{C}(\mathcal{P},m,c)$ is to derive the minimax testing radius $c$ such that the conditionally dependent and independent instances can be distinguished accurately. Crucially, the hypotheses $\mathcal{H}_{0}\text{ and }\mathcal{H}_{1}$, together with the dependence measure $m$, can be used to characterize the $c$-strong tree-faithfulness. We continue to introduce the poly-forest learning problem using this set-up.
Definition 4 (Poly-forest learning problem).
The poly-forest learning problem $\mathcal{F}(\mathcal{P},m,c)$ is to learn the Markov equivalence class of a poly-forest $G\in\mathcal{T}$ given i.i.d. samples from a distribution $p$ that is Markov and $c$-strong tree-faithful to $G$ with respect to the dependence measure $m$, and for any triplet of nodes $(j,k,\ell)$, the marginal $p(X_{j},X_{k},X_{\ell})$ comes from either $\mathcal{H}_{0}$ or $\mathcal{H}_{1}$. The optimal sample complexity of $\mathcal{F}(\mathcal{P},m,c)$ refers to the sufficient and necessary conditions on sample size $n$ in terms of the number of nodes $d$ and the strong tree-faithfulness parameter $c$, such that for some estimator $\widehat{G}$ and some constant $\delta\in(0,1)$, the following holds:
$\sup_{P}\mathbb{P}(\widehat{G}\neq\overline{G})\leq\delta\,.$
The poly-forest model is defined by requiring the marginal distributions of all node triplets to fall into the hypothesis classes $\mathcal{H}_{0}$ or $\mathcal{H}_{1}$, which follow from $\mathcal{P}$ and $m$. This requirement is minimal as it essentially repeats the strong tree-faithfulness while defining $p$ to share the distributional properties of the distributions considered in the CI testing problem, e.g. Gaussianity or smoothness. We focus on the optimal sample complexity of recovering the CPDAG for the poly-forest in $\mathcal{F}(\mathcal{P},m,c)$.
To provide an estimator that achieves the optimal sample complexity for $\mathcal{F}(\mathcal{P},m,c)$, we adopt the PC-tree algorithm introduced in Wang et al. (2024). PC-tree determines the edges between any two nodes by testing their (conditional) independence given one other node, and is shown to consistently learn the Gaussian poly-tree models. The algorithm takes a valid CI test as input, is efficient and runs in polynomial time in number of nodes $d$, and is readily extended for poly-forests for general distributions. We detail the specifics of PC-tree algorithm in Appendix A.
3 Equivalence between CI testing and structure learning
We present our main result below, establishing a fundamental connection between CI testing $\mathcal{C}(\mathcal{P},m,c)$ and structure learning $\mathcal{F}(\mathcal{P},m,c)$.
Theorem 3.1.
Given a conditional independence testing problem $\mathcal{C}(\mathcal{P},m,c)$ with an optimal test $\psi$ achieving the minimax testing radius $c\asymp n^{-1/\alpha}$, if there exist hard instances $p_{0}\in\mathcal{H}_{0}$ and $p_{1}\in\mathcal{H}_{1}$ that are Markov and $c$-strong tree-faithful, then the optimal sample complexity of learning $\mathcal{F}(\mathcal{P},m,c)$ is
$n\asymp\frac{\log d}{c^{\alpha}}\,,$ (1)
which is achieved by PC-tree with $\psi$.
Theorem 3.1 establishes a fundamental statistical reduction between the two problems: the optimality of a CI testing problem implies the optimality of corresponding structure learning for poly-forests. In particular, the inherent difficulty in CI testing directly translates to the difficulty in poly-forest learning, and once the minimax solution (an optimal CI test $\psi$) is obtained, it can be directly plug-in for learning the poly-forest. A full technical statement of this result is deferred to Theorem C.1 in Appendix C, where also additional discussion and the proof can be found.
The condition in Theorem 3.1 looks for two hard instances $p_{0}\in\mathcal{H}_{0}$ and $p_{1}\in\mathcal{H}_{1}$ that are difficult to distinguish in the sense of small KL-divergence, see (C2) for the detailed requirement. We stress that this condition is imposed merely on triplet of variables (cf. Definition 3), and is a standard step to establish lower bounds when studying the minimax optimality for CI testing. Additional requirements on the graphical properties, i.e. Markov and strong tree-faithfulness, of these instances are introduced to ensure their validity in the context of poly-forest learning. As will be shown in Section 5, this technical requirement sometimes necessitates minor modifications on existing constructions in the literature of CI testing.
In the optimal sample complexity of structure learning (1), the effect of dimensionality comes in as a factor of $\log d$, and the exponent $\alpha$ on dependence of signal draws connection between CI testing and structure learning problems, as it directly translates from the optimal testing radius to the optimal sample complexity. For parametric distributions, we typically have $\alpha=2$, corresponding to the parametric rate. While for nonparametric distributions, it is often the case that $\alpha>2$ and depends on the smoothness conditions. We provide examples on both cases in Sections 4-5.
To apply Theorem 3.1 on concrete problems to obtain optimality, one needs to specify the distributional assumptions in the CI testing problem, design an optimal CI test, and find proper instances from the (in)dependent hypotheses that are close enough in KL-divergence. In the remainder of this paper, we show how to apply Theorem 3.1 to Bernoulli, Gaussian, and nonparametric continuous distributions. In practice, given a powerful test $\psi$ tailored to the models satisfying certain distributional assumptions, one can directly integrate it into PC-tree algorithm for efficient poly-forest learning. We demonstrate this across the three distributional settings in the experiments in Section 6.
4 Applications to Bernoulli and Gaussian distributions
In this and the following sections, we illustrate by examples the broad applicability of the main optimality result (Theorem 3.1) through a series of representative distributions. For each example, we specify the associated model class, provide a valid CI test, and give hard instances that satisfy the condition of Theorem 3.1. For the sake of space, we state and discuss the key conclusions in the main paper and defer full details of the model class descriptions to Appendix B.
4.1 Bernoulli distribution
We start with Bernoulli distribution. As mentioned, it suffices to define the CI testing problem to set the stage. We consider all multivariate Bernoulli distributions with dimension being three, i.e. $\mathbb{P}(X=x,Y=y,Z=z)=p(x,y,z),\forall(x,y,z)\in\{0,1\}^{3}$ and $\sum_{x,y,z}p(x,y,z)=1$. Formal details are deferred to Appendix B.1. We measure dependence using
$m^{B}(X,Y\,|\,Z)=\sum_{z}p(z)\sum_{x,y}\Big{|}p(x,y\,|\,z)-p(x\,|\,z)p(y\,|\,z)\Big{|}\,,$
which is the total variation distance with the product of its marginals and is a valid dependence measure. These elements lead to CI testing problem $\mathcal{C}^{B}$ and the poly-forest learning problem $\mathcal{F}^{B}$, where $B$ denotes “Bernoulli”. $\mathcal{F}^{B}$ effectively contains all multivariate Bernoulli distributions with dimension $d$ and Markov and $c$-strong tree-faithful to some poly-forest $G\in\mathcal{T}$:
$(X_{1},X_{2}\ldots,X_{d})\sim\mathbb{P}(X_{1}=x_{1},X_{2}=x_{2},\ldots,X_{d}=x_{d}),\quad(x_{1},x_{2}\ldots,x_{d})\in\{0,1\}^{d}\,.$
Now we proceed to construct a test based on thresholding the estimation of the dependence measure by the sample counterpart, inspired by the classical $\chi^{2}$ independence test:
$\widehat{m}^{B}(X,Y\,|\,Z)$ $=\sum_{z}\widehat{p}(z)\sum_{x,y}\Big{|}\widehat{p}(x,y\,|\,z)-\widehat{p}(x\,|\,z)\widehat{p}(y\,|\,z)\Big{|}$
$\psi^{B}$ $=\mathbbm{1}\{\widehat{m}^{B}(X,Y\,|\,Z)\geq c/2\}$ (2)
where $\widehat{p}(x)=\sum_{i}\mathbbm{1}\{X_{i}=x\}/n$ for $x\in\{0,1\}$ and other estimates are analogously defined.
For the hard instance constructions, we consider $p_{0}^{B}(X,Y,Z)$ under which $X,Y,Z$ are independent $\text{Bern}(\frac{1}{2})$ random variables, and the alternative distribution $p^{B}_{1}(X,Y,Z)$ is given as follows:
$Z\sim\text{Bern}\Big{(}\frac{1}{2}\Big{)},\quad X\,|\,Z\sim\begin{cases}\text{Bern}(\frac{1}{2}+c)&Z=1\\ \text{Bern}(\frac{1}{2}-c)&Z=0\end{cases},\quad Y\,|\,X\sim\begin{cases}\text{Bern}(\frac{1}{2}+c)&X=1\\ \text{Bern}(\frac{1}{2}-c)&X=0\end{cases}\,.$
It is easy to see that $p_{0}^{B}$ is constructed to be Markov and tree-faithful to an empty graph, while $p_{1}^{B}$ is constructed to follow a three-node chain graph $Z\to X\to Y$, both of which are poly-forests.
Then the following theorem checks the validity of the test and construction, and concludes the optimality of learning Bernoulli poly-forest via Theorem 3.1. See proof in Appendix D.
Theorem 4.1.
$\psi^{B}$ is optimal for $\mathcal{C}^{B}$ with optimal testing radius being $c\asymp n^{-1/2}$, and the optimal sample complexity of learning $\mathcal{F}^{B}$ is
$n\asymp\frac{\log d}{c^{2}}\,,$
which is achieved by PC-tree with $\psi^{B}$.
4.2 Gaussian distribution
We next consider Gaussian distribution as another parametric example to illustrate the main result. Since this application will recover the established optimality in Wang et al. (2024), we only collect the key elements and defer most details to Appendix B.2. We consider all multivariate Gaussian distributions with dimension being three and measure dependence using the partial correlation coefficient (cf. (3)). The Gaussian CI testing problem $\mathcal{C}^{G}$ and the Gaussian poly-forest learning problem $\mathcal{F}^{G}$ are defined, where $G$ represents “Gaussian”.
We employ a valid CI test $\psi^{G}$ similar to $\psi^{B}$ by thresholding the sample partial correlation (cf. (4)). For the hard instances, let $p_{0}^{G}$ be $\mathcal{N}(\mathbf{0}_{3},I_{3})$ and $p_{1}^{G}$ be generated as
$Z\sim\mathcal{N}(0,1),\quad X=\beta Z+\mathcal{N}(0,1-\beta^{2}),\quad Y=\beta X+\mathcal{N}(0,1-\beta^{2})\,,$
where $\beta=2c$. Likewise in the Bernoulli case, $p_{0}^{G}$ is constructed to be Markov and tree-faithful to an empty graph, and $p_{1}^{G}$ is constructed for the chain graph $Z\to X\to Y$. It remains to check the validity of $\psi^{G}$, $p_{0}^{G}$ and $p_{1}^{G}$. Applying Theorem 3.1, we have the optimality of learning Gaussian poly-forest (proof is in Appendix E):
Theorem 4.2.
$\psi^{G}$ is optimal for $\mathcal{C}^{G}$ with optimal testing radius being $c\asymp n^{-1/2}$, and the optimal sample complexity of learning $\mathcal{F}^{G}$ is $n\asymp\log d/c^{2}$, which is achieved by PC-tree with $\psi^{G}$.
5 Application to nonparametric models
For both Bernoulli and Gaussian distributions, the optimal sample complexity for learning poly-forest is $\Theta(\log d/c^{2})$, i.e. $\alpha=2$ in Theorem 3.1. This result primarily arises from the parametric nature of these distributions. To explore the implications beyond the parametric setting, we now shift our focus to a nonparametric continuous distribution. This allows us to examine how the nonparametric nature influences the value of $\alpha$, leading to a distinct theoretical outcome.
We adopt the framework in Neykov et al. (2021) and provide the essential elements in Appendix B.3, see details therein. As alluded to in Example 1, we consider all continuous distributions over $[0,1]^{3}$, which admit continuous densities $p(X,Y,Z)$. We measure the dependence using
$m^{NP}(X,Y\,|\,Z)=\inf_{q\in\mathcal{Q}}\|p-q\|_{1}\,,$
[…]
Now we proceed to verify the applicability of Theorem 3.1 for this nonparametric setting. We use the minimax optimal CI test introduced in Section 5.3 of Neykov et al. (2021), denoted as $\psi^{NP}$, which is based on classic U-statistics to measure the (conditional) dependence between the $X$ and $Y$. $\psi^{NP}$ is a valid choice and satisfies the type I and II error controls with $\alpha=\frac{5s+2}{2s}$ (see Theorem 5.6 therein).
For the hard instances, we design the constructions, denoted as $p_{0}^{NP}$ and $p_{1}^{NP}$, based on a modification of the ones used for proving lower bounds for nonparametric CI testing in Neykov et al. (2021). Although the original constructions of $p_{0}$ and $p_{1}$ are close in KL divergence, the issue lies in the unfaithfulness of the distribution, thus they cannot be directly applied for poly-forest learning setting. Specifically, the original construction of the alternative $p_{1}$ only characterizes the conditional dependence but accidentally leads to marginal independence between variables, which cannot be faithful to any poly-forest (over triplet of nodes).
We modify the construction by adding back the marginal dependence with extra complexity in the analysis. Specifically, let $p_{0}^{NP}$ be independent uniform distributions $Unif^{3}[0,1]$, which is Markov to the empty graph. For $p_{1}^{NP}$, we consider a $V$-structure $X\rightarrow Z\leftarrow Y$, which is a three-node poly-forest. Under this graph, a faithful distribution is supposed to have $X\perp\!\!\!\perp Y$ while $X\not\perp\!\!\!\perp Y\,|\,Z$ and $X\not\perp\!\!\!\perp Z,Y\not\perp\!\!\!\perp Z$. Thus $p_{1}^{NP}$ is specified as follows: $X,Y\sim Unif[0,1]$ and $p_{Z\,|\,X,Y}$ is given by a perturbation to the uniform depending on $X$ and $Y$:
$p(Z\,|\,X,Y)=1+\gamma_{\Delta}(X,Y)\eta_{\nu}(Z)\,,$
for some functions $\gamma_{\Delta}$ and $\eta_{\nu}$ governing the perturbation, which are specified in the proof. This construction is in the same form as the one in Neykov et al. (2021). However, the functions are constructed such that the $X$ (or $Y$) and $Z$ are marginally dependent, thus faithfulness is satisfied. See more details about these constructions in the proof (Appendix F). Applying Theorem 3.1, we have the optimality of learning nonparametric continuous poly-forest.
Theorem 5.1.
$\psi^{NP}$ is optimal for $\mathcal{C}^{NP}$ with optimal testing radius being $c\asymp n^{-2s/(5s+2)}$, and the optimal sample complexity of learning $\mathcal{F}^{NP}$ is
$n\asymp\frac{\log d}{c^{\frac{5s+2}{2s}}}\,,$
which is achieved by PC-tree with $\psi^{NP}$.
Theorem 5.1 highlights the larger sample complexity of nonparametric setting compared to the parametric one for poly-forest learning. The difference lies in the dependence on the strong tree-faithfulness parameter $c$, where the resulting exponent $\alpha=\frac{5s+2}{2s}>2$ comes directly from the intrinsic hardness of nonparametric CI testing. Moreover, the modification for $p_{0}^{NP}$ and $p_{1}^{NP}$ illustrates the additional complexity in the analysis to ensure faithfulness.
6 Experiments
We conducted experiments to validate our theoretical findings on CI testing and structure learning with PC-tree algorithm (detailed in Algorithm 1). Specifically, we conducted a series of simulation studies for structure learning in Bernoulli, Gaussian, and nonparametric continuous distributions, corresponding to the examples provided in Sections 4-5. We simulate poly-forest models for each distribution setting and apply PC-tree algorithm equipped with the CI tests $\psi^{B}$, $\psi^{G}$, and $\psi^{NP}$ specified in Sections 4-5 to estimate the graph structure.
Figure 1: Structure Hamming distance (SHD) vs. sample size for poly-forest learning for Bernoulli, Gaussian, and nonparametric continuous distributions over varying number of nodes (indicated by colors). Error bars represent standard deviations. SHD consistently decreases toward zero as sample size increases across all experimental settings.
In Figure 1, we present the structure Hamming distance (SHD) between the estimated graph and the ground truth against sample size $n$ for various settings of number of nodes $d=\{20,40,60,80,100\}$. Each setting is averaged over $N=20$ replications. The results show as the sample size increases, SHD consistently decreases for graphs with 20 to 100 nodes. A lower SHD means the output graph is closer to the truth, thus more accurate structure learning. The effect of dimensionality is illustrated by lines with different colors. Overall, our simulations demonstrate that combining a powerful conditional independence test $\psi$ with PC-tree algorithm allows for consistent and accurate poly-forest learning. Full experiment details on implementations and how the data are simulated, and additional experiment results are postponed to Appendix H.
7 Conclusion
In this paper, we study the minimax optimality of structure learning and conditional independence testing. We establish a connection between CI testing and structure learning by showing the optimality conditions for CI testing translate directly into the optimality of poly-forest learning, which can be achieved by an efficient constraint-based algorithm with the optimal CI test as input. The generic theoretical results are demonstrated using three representative distribution families. This finding highlights the central role of CI testing in structure learning in the view of statistical sample efficiency.
One interesting direction for future research is to extend the results beyond poly-forests to general DAGs. In such settings, constraint-based methods — such as the PC algorithm — are applicable and rely solely on CI testing to operate. We anticipate that a similar statistical connection between CI testing and structure learning can be established for general DAGs, though additional structural parameters, such as the maximum in-degree, are likely to play a key role in characterizing the corresponding optimal sample complexity.
Algorithm 1 PC-Tree algorithm
Input: $n$ i.i.d. samples $\{X^{(i)}_{1},\ldots,X^{(i)}_{d}\}_{i=1}^{n}$, CI test $\psi$ as function of data;
  1. 1.
    Let $\widehat{E}=\emptyset$.
  2. 2.
    For each pair $(j,k)$, $0\leq j<k\leq d$:
    1. (a)
      For all $\ell\in[d]\cup\{\emptyset\}\setminus\{j,k\}$:
      1. i.
        Test $H_{0}:X_{j}\perp\!\!\!\perp X_{k}\,|\,X_{\ell}$ vs. $H_{1}:X_{j}\not\perp\!\!\!\perp X_{k}\,|\,X_{\ell}$ using $\psi$, store the results.
    2. (b)
      If all tests reject, then $\widehat{E}\leftarrow\widehat{E}\cup\{j-k\}$.
    3. (c)
      Else (if some test accepts), let $S(j,k)=\{\ell\in[d]\cup\{\emptyset\}\setminus\{j,k\}:X_{j}\perp\!\!\!\perp X_{k}\,|\,X_{\ell}\}$.
Output: $\widehat{G}=([d],\widehat{E})$, separation set $S$.
Algorithm 2 Orient algorithm
Input: Skeleton $\widehat{G}$, separation sets $S$
Output: CPDAG $\widehat{\overline{G}}$.
  1. 1.
    For all pairs of nonadjacent nodes $j,k$ with common neighbour $\ell$:
    1. (a)
      If $\ell\not\in S(j,k)$, then directize $j-\ell-k$ in $\widehat{G}$ by $j\rightarrow\ell\leftarrow k$
  2. 2.
    In the resulting PDAG $\widehat{G}$, orient as many as possible undirected edges by applying following rules:
    • R1 Orient $k-\ell$ into $k\rightarrow\ell$ whenever there is an arrow $j\rightarrow k$ such that $j$ and $\ell$ are not adjacent
    • R2 Orient $j-k$ into $j\rightarrow k$ whenever there is a chain $j\rightarrow\ell\rightarrow k$
    • R3 Orient $j-k$ into $j\rightarrow k$ whenever there are two chains $j-\ell\rightarrow k$ and $j-i\rightarrow k$ such that $\ell$ and $i$ are not adjacent
    • R4 Orient $j-k$ into $j\rightarrow k$ whenever there are two chains $j-\ell\rightarrow i$ and $\ell-i\rightarrow k$ such that $\ell$ and $i$ are not adjacent
  3. 3.
    Return $\widehat{G}$ as $\widehat{\overline{G}}$.
Appendix A Details of PC-tree algorithm
Introduced in Wang et al. (2024), PC-tree algorithm (Algorithm 1) is a modification to the classical PC algorithm and tailored for learning poly-forests/trees. Generically, it takes observational data along with a valid CI test as input, and outputs the skeleton with a separation set. The estimated skeleton will be further oriented by rules specified in Algorithm 2 using the separation set.
PC-tree conducts CI tests to determine the presence of edge between any two nodes. To achieve this, it only tests marginal independence and conditional independence given only one other node, rather than considering all possible conditioning sets as in the classical PC algorithm. Therefore, PC-tree only invokes
$\binom{d}{2}\times\Big{[}1+(d-2)\Big{]}\asymp d^{3}$
many times of CI test $\psi$, thereby is efficient. Since poly-forests are effectively concatenation of poly-trees, PC-tree is also consistent for learning poly-forest structures. As long as the input CI test $\psi$ is valid and well controls the type-I and type-II errors, the consistency of the algorithm applies to general distributions.
Appendix B Details of applications
In this appendix, we detail the formal definitions of the model class considered in each application (Sections 4-5). Specifically, we specify $\mathcal{P}$ and $m$, thereby $\mathcal{H}_{1}$ and $\mathcal{H}_{0}$ will follow, for the exampled distributions.
B.1 Bernoulli distribution (Section 4.1)
We start by specifying the model class $\mathcal{P}$. Consider all multivariate Bernoulli distributions with dimension being three. They are parametrized by joint probability mass function $p(x,y,z)$:
[…]
$\sum_{x,y,z}p(x,y,z)=1\bigg{\}}\,.$
We measure dependence using the total variation distance to the product of its marginals:
$m^{B}(X,Y\,|\,Z)=\sum_{z}p(z)\sum_{x,y}|p(x,y\,|\,z)-p(x\,|\,z)p(y\,|\,z)|\,.$
Then $\mathcal{H}_{0}^{B}$ and $\mathcal{H}_{1}^{B}$ contain all these Bernoulli distributions that are conditionally independent and dependent respectively:
$\mathcal{H}_{0}^{B}=\{p\in\mathcal{P}^{B}:X\perp\!\!\!\perp Y\,|\,Z\}\,,\quad\mathcal{H}_{1}^{B}=\{p\in\mathcal{P}^{B}:m^{B}(X,Y\,|\,Z)\geq c\}$
Having defined the elements of CI testing problem $\mathcal{C}^{B}:=\mathcal{C}(\mathcal{P}^{B},m^{B},c)$, the Bernoulli poly-forest learning problem $\mathcal{F}^{B}:=\mathcal{F}(\mathcal{P}^{B},m^{B},c)$ contains all multivariate Bernoulli distributions with dimension $d$ and Markov and $c$-strong tree-faithful to some poly-forest $G\in\mathcal{T}$:
$(X_{1},X_{2}\ldots,X_{d})\sim\mathbb{P}(X_{1}=x_{1},X_{2}=x_{2},\ldots,X_{d}=x_{d}),\quad(x_{1},x_{2}\ldots,x_{d})\in\{0,1\}^{d}\,.$
B.2 Gaussian distribution (Section 4.2)
Consider all multivariate Gaussian distributions with dimension being three:
$\mathcal{P}^{G}=\bigg{\{}p(X,Y,Z):(X,Y,Z)\sim\mathcal{N}(\mathbf{0}_{3},\Sigma),\Sigma\in\mathbb{S}_{++}^{3}\bigg{\}}\,.$
We measure dependence using the partial correlation:
$m^{G}(X,Y\,|\,Z)=\frac{|\operatorname{cov}(X,Y\,|\,Z)|}{\sqrt{\operatorname{var}(X\,|\,Z)\operatorname{var}(Y\,|\,Z)}}\,.$ (3)
Then $\mathcal{H}_{0}^{G}$ and $\mathcal{H}_{1}^{G}$ contain all these Gaussian distributions that are conditionally independent and dependent respectively.
$\mathcal{H}_{0}^{G}=\{p\in\mathcal{P}^{G}:X\perp\!\!\!\perp Y\,|\,Z\}\,,\quad\mathcal{H}_{1}^{G}=\{p\in\mathcal{P}^{G}:m^{G}(X,Y\,|\,Z)\geq c\}$
Consequently, the Gaussian CI testing problem $\mathcal{C}^{G}:=\mathcal{C}(\mathcal{P}^{G},m^{G},c)$ is defined. Meanwhile, it gives the Gaussian poly-forest learning problem $\mathcal{F}^{G}:=\mathcal{F}(\mathcal{P}^{G},m^{G},c)$, which include all multivariate Gaussian distributions with dimension $d$ and Markov and $c$-strong tree-faithful to some poly-forest $G\in\mathcal{T}$:
$(X_{1},X_{2}\ldots,X_{d})\sim\mathcal{N}(\mathbf{0}_{d},\Sigma),\Sigma\in\mathbb{S}_{++}^{d}\,.$
Similar to $\psi^{B}$, we construct a test by thresholding the sample partial correlation:
[…]
where $\widehat{\Sigma}=\frac{1}{n}\sum_{i=1}^{n}X^{(i)}{X^{(i)}}^{\top}$ is the sample covariance matrix.
B.3 Nonparametric continuous distribution (Section 5)
We follow the CI testing framework in Neykov et al. (2021). Consider all continuous distributions over $[0,1]^{3}$ that admit continuous densities $p(X,Y,Z)$. In addition, we impose following two smoothness condition on the densities.
Definition 5 (Lipschitzness).
For some constant $L_{1}>0$,
  • if $X\perp\!\!\!\perp Y\,|\,Z$, then $\|p_{X\,|\,Z=z}-p_{X\,|\,Z=z}\|_{1}\leq L_{1}|z-z^{\prime}|$ and $\|p_{Y\,|\,Z=z}-p_{Y\,|\,Z=z^{\prime}}\|_{1}\leq L_{1}|z-z^{\prime}|$.
  • if $X\not\perp\!\!\!\perp Y\,|\,Z$, then $\|p_{X,Y\,|\,Z=z}-p_{X,Y\,|\,Z=z^{\prime}}\|_{1}\leq L_{1}|z-z^{\prime}|$.
Definition 6 ($s$-Hölder smoothness).
For some $s>0$, let $\lfloor s\rfloor$ be the maximum integer smaller than $s$. For some constant $L_{2}>0$ and all $x,x^{\prime},y,y^{\prime},z\in[0,1]$,
$\sup_{t\leq\lfloor s\rfloor}\bigg{|}\frac{\partial^{t}}{\partial x^{t}}\frac{\partial^{\lfloor s\rfloor-t}}{\partial y^{\lfloor s\rfloor-t}}p(x,y\,|\,z)-\frac{\partial^{t}}{\partial x^{t}}\frac{\partial^{\lfloor s\rfloor-t}}{\partial y^{\lfloor s\rfloor-t}}p(x^{\prime},y^{\prime}\,|\,z)\bigg{|}\leq L_{2}\sqrt{(x-x^{\prime})^{2}+(y-y^{\prime})^{2}}^{s-\lfloor s\rfloor}$
and
$\sup_{t\leq\lfloor s\rfloor}\bigg{|}\frac{\partial^{t}}{\partial x^{t}}\frac{\partial^{\lfloor s\rfloor-t}}{\partial y^{\lfloor s\rfloor-t}}p(x,y\,|\,z)\bigg{|}\leq L_{2}\,.$
The Lipschitzness is used to characterize the smoothness with respect to the conditioning variable $Z$, while the Hölder smoothness is required for the conditional densities with the conditioning variable fixed, and particularly for the (conditionally) dependent variable pairs. Denote this class of distributions as $\mathcal{P}^{NP}$. As specified in the main paper, we measure the dependence using
$m^{NP}(X,Y\,|\,Z)=\inf_{q\in\mathcal{Q}}\|p-q\|_{1}$
which is the distance to the closest (conditionally) independent distributions, and should be large enough such that the dependent distributions can be distinguished from the independent ones (Shah and Peters, 2020). Therefore, the CI testing class is defined $\mathcal{C}^{NP}:=\mathcal{C}(\mathcal{P}^{NP},m^{NP},c)$, with the null and alternative hypotheses being
$\mathcal{H}_{0}^{NP}$ =$=\{p\in\mathcal{P}^{NP}:X\perp\!\!\!\perp Y\,|\,Z\text{ and }\text{satisfies Lipschitzness}\}$
[…]
Correspondingly, the nonparametric poly-forest learning problem is defined, which includes all continuous distributions supported on $[0,1]^{d}$ and Markov and $c$-strong tree-faithful to some poly-forest $G\in\mathcal{T}$, denoted as $\mathcal{F}^{NP}:=\mathcal{F}(\mathcal{P}^{NP},m^{NP},c)$.
Appendix C Proof of Theorem 3.1
We provide the full version of Theorem 3.1 below, followed by the discussion and proof.
Theorem C.1.
Given a conditional independence testing problem $\mathcal{C}(\mathcal{P},m,c)$, if for some constants $A$ and $A^{\prime}$,
  1. (C1)
    There exists a test $\psi$ such that when $c\geq A\times n^{-1/\alpha}$, it holds that
    $\sup_{p\in\mathcal{H}_{0}}\mathbb{E}_{p}[\psi]+\sup_{p\in\mathcal{H}_{1}}\mathbb{E}_{p}[1-\psi]\leq\frac{1}{2}\,;$
  2. (C2)
    There exist $p_{0}\in\mathcal{H}_{0}$ and $p_{1}\in\mathcal{H}_{1}$ such that $\mathbf{KL}(p_{1}\|p_{0})\leq A^{\prime}\times c^{\alpha}$;
then the minimax testing radius of $\mathcal{C}(\mathcal{P},m,c)$ is $c\asymp n^{-1/\alpha}$, and $\psi$ is optimal for $\mathcal{C}(\mathcal{P},m,c)$. In addition, if
  1. (C3)
    $p_{0}$ and $p_{1}$ are Markov and $c$-strong tree-faithful to some poly-forests of three nodes;
then the optimal sample complexity of learning $\mathcal{F}(\mathcal{P},m,c)$ is
$n\asymp\frac{\log d}{c^{\alpha}}\,,$
which is achieved by PC-tree with $\psi$.
The first two conditions assumed in Theorem C.1 are standard steps to study minimax optimality, corresponding to upper and lower bounds of testing radius. It is typical in literature to verify these two conditions when deriving the optimal testing radius:
  • (C1) requires finding a powerful test to distinguish independent and dependent instances with sufficiently large signal.
  • (C2) looks for two instances from $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ that are hard to distinguished. Typically, $p_{0}$ is a “flat” distribution, e.g. uniform distribution or $Bern(0.5)$, while $p_{1}$ is constructed to be a perturbation to the flat distribution with sufficiently large signal $c$ but small deviation to being independent.
  • With above two conditions, to conclude optimal poly-forest learning, (C3) requires the instances to be generated by poly-forest models, ensuring their validity in the context of structure learning.
Proof.
Under the given conditions, we prove the optimalities for CI testing and poly-forest learning.
Optimal CI testing
The existence of test $\psi$ implies a upper bound of testing radius $c\lesssim n^{-1/\alpha}$. The lower bound is given by Le Cam’s two point method: by the existence $p_{0}\in\mathcal{H}_{0},p_{1}\in\mathcal{H}_{1}$, we have
$\inf_{\psi}\bigg{\{}\sup_{p\in\mathcal{H}_{0}}\mathbb{E}_{p}[\psi]+\sup_{p\in\mathcal{H}_{1}}\mathbb{E}_{p}[1-\psi]\bigg{\}}\geq\frac{1}{2}\Big{(}1-n\mathbf{TV}(p_{1}\|p_{0})\Big{)}\,,$
and $\mathbf{TV}(p_{1}\|p_{0})\leq\mathbf{KL}(p_{1}\|p_{0})\leq A^{\prime}\times c^{\alpha}$. This requires the testing radius $nc^{\alpha}\gtrsim 1$, which yields $c\gtrsim n^{-1/\alpha}$.
Optimal poly-forest learning
Regarding the upper bound, we start with applying median trick (Motwani, 1995) to obtain an exponential decay of error probability. The idea is to divide the full sample into $K$ folds and apply the statistical test on each sub-sample. Take the majority vote of all these tests as the final output. Then the final output makes mistake only when half of the tests make mistake, whose probability can easily computed and bounded in an exponential way and serves our goal when $K$ is appropriately chosen.
Proposition C.2.
Suppose there exists a statistical test $\psi$ such that if the sample size $n\gtrsim N$, then
$\mathbb{P}(\psi\text{ is incorrect})\leq 1/2\,,$
then there exists a CI test $\psi^{\prime}$ such that under the same condition,
$\mathbb{P}(\psi^{\prime}\text{ is incorrect})\leq\exp\Big{(}-C_{0}n/N\Big{)}\,.$
Applying Proposition C.2, we can derive an error probability bound for $\psi$ with modification. For simplicity, we will stick with $\psi$ with a little abuse of notation. Assuming $n\gtrsim c^{-\alpha}$, we have for some constant $C_{0}$,
$\mathbb{P}(\psi\text{ is incorrect})\leq\exp\Big{(}-C_{0}nc^{\alpha}\Big{)}\,.$
With this error bound in hand, the proof proceeds by following the proof of Theorem 4.3 in Wang et al. (2024) with minor modifications: 1) replacing poly-tree with poly-forest; 2) replacing correlation CI testing with $\psi$. This completes the proof of upper bound, and shows PC-tree with $\psi$ as CI tester achieves the upper bound.
Regarding the lower bound, without loss of generality, we assume $d$ divides $3$, otherwise proceed by setting the remaining one or two nodes to be isolated and follow $p_{0}(X)$. Group all the nodes into $d^{\prime}=d/3$ clusters:
$\{1,2,3\},\{4,5,6\},\cdots,\{3k+1,3k+2,3k+3\},\cdots,\{3d^{\prime}+1,3d^{\prime}+2,3d^{\prime}+3\}\,.$
We are going to use the $p_{1},p_{0}$ to parametrize the construction. Since $p_{1}$ and $p_{0}$ give different conditional independence statements, they are Markov and $c$-strong tree-faithful to two different poly-forests, which are denoted as $T_{0},T_{1}$.
Firstly, construct graph $G_{0}$ by stacking $d^{\prime}$ many $T_{0}$’s together. Then for each $k\geq 1$, construct graph $G_{k}$ by stacking $d^{\prime}-1$ many $T_{0}$’s and one $T_{1}$ together, and the subgraph $T_{1}$ is imposed on nodes $\{3k+1,3k+2,3k+3\}$, while $T_{0}$ is imposed on all remaining triplets. Under this construction, $\{G_{k}\}_{k\geq 0}$ are poly-forests with (at least) $d^{\prime}$ many disconnected subgraphs, and are distinct with each other.
Now we consider distributions for each $k\geq 0$. Let $P_{0}=p_{0}^{\otimes d^{\prime}}$ and
$P_{k}(X)=p_{0}^{\otimes d^{\prime}-1}\times P_{k}(X_{3k+1},X_{3k+2},X_{3k+3})\quad\text{ with }\quad P_{k}(X_{3k+1},X_{3k+2},X_{3k+3})=p_{1}\,.$
Since $p_{0}$ and $p_{1}$ are Markov and $c$-strong tree-faithful to $T_{0}$ and $T_{1}$ respectively, $P_{k}$ is Markov and $c$-strong tree-faithful to $G_{k}$ for all $k\geq 0$. Therefore, $\{P_{k}\}_{k\geq 0}\in\mathcal{F}(\mathcal{P},m,c)$.
We also going to apply Tsybakov’s method (Corollary G.4). Firstly, we have the size of the construction to be lower bounded
$\log M=\log(d^{\prime}+1)\geq\frac{1}{2}\log d\,.$
Then we upper bound the KL divergence between $P_{k}$ and $P_{0}$ for all $k\geq 1$:
$\mathbf{KL}(P_{k}\|P_{0})=\mathbb{E}_{P_{k}}\log\frac{P_{k}}{P_{0}}=\mathbb{E}_{P_{k}}\log\frac{p_{1}}{p_{0}}=\mathbf{KL}(p_{1}\|p_{0})\leq A^{\prime}\times c^{\alpha}\,.$
Invoking Corollary G.4 completes the proof. ∎
Proof of Proposition C.2.
Divide the full sample into $K$ folds, which we will specify later, and apply $\psi_{0}$ for each of the sub-sample to get outputs $\psi^{(1)},\ldots,\psi^{(K)}$. Let
$\psi=\operatorname*{arg\,max}_{t\in\{0,1\}}\sum_{k=1}^{K}\mathbbm{1}\{\psi^{(k)}=t\}$
be the majority vote of all the tests as final output. Therefore, the incorrectness of $\psi$ implies at least half of the sub-tests make mistakes. Suppose that $n/K\gtrsim N$, by the guarantee of $\psi_{0}$, we have
$\mathbb{P}(\psi\text{ is incorrect})$ $\leq\mathbb{P}\Big{(}\text{half of }\{\psi^{(k)}\}_{k=1}^{K}\text{ is incorrect}\Big{)}$
$\leq\frac{1}{2^{K/2}}=\exp\Big{(}-\frac{\log 2}{2}\times K\Big{)}\,.$
Now set $K=C^{\prime}n/N$ for constant $C^{\prime}$ large enough such that $n/K\gtrsim N$ is satisfied, let $C_{0}=C^{\prime}(\log 2)/2$, we obtain
$\mathbb{P}(\psi\text{ is incorrect})\leq\exp\Big{(}-C_{0}n/N\Big{)}$
as desired. ∎
Appendix D Proof of Theorem 4.1 (Bernoulli distribution)
We start by showing the upper bound (C1) of $\psi^{B}$, then proceed to verify the validity of $p_{0}^{B}$ and $p_{1}^{B}$ and show they satisfy (C2) and (C3).
D.1 Proof of upper bound for Bernoulli distribution
Proof.
We directly show $\psi^{B}$ satisfies an exponential bound. Notice that by construction, the concentration of $\widehat{m}$ implies the correctness of testing:
$|\widehat{m}^{B}-m^{B}|\leq c/2\implies\psi^{B}\text{ is correct.}$.
We aim to show $|\widehat{m}^{B}-m^{B}|\leq c/2$ holds with high probability. To achieve this, we employee the result below:
Lemma D.1 (Devroye (1983), Lemma 3).
Let $(X_{1},X_{2},\ldots,X_{k})$ be a multinomial $(n,p_{1},p_{2},\ldots,p_{k})$ random vector. Let $\widehat{p}=(X_{1},X_{2},\ldots,X_{k})/n$ For all $\epsilon\in(0,1)$ and all $k$ satisfying $k/n\leq\epsilon^{2}/20$, we have
$\mathbb{P}(\|\widehat{p}-p\|_{1}>\epsilon)\leq 3\exp(-n\epsilon^{2}/25)\,.$
Applied to our setup, we consider the concentration for the joint distribution $p_{XYZ}$, which can be viewed as a multinomial distribution with dimension $k=2^{3}=8$. Therefore, for $\epsilon$ such that $n\geq 160/\epsilon^{2}$, with probability at least $1-3\exp(-n\epsilon^{2}/25)$, we have
$\|p_{XYZ}-\widehat{p}_{XYZ}\|_{1}:=\sum_{xyz}|p(x,y,z)-\widehat{p}(x,y,z)|\leq\epsilon\,,$
which also implies the concentration of $p_{WZ}$ ($W\in\{X,Y\}$) and $p_{Z}$. To see this, take $W=X$ for example:
$\|\widehat{p}_{XZ}-p_{XZ}\|_{1}$ $=\sum_{xz}|\widehat{p}(x,z)-p(x,z)|=\sum_{xz}|\sum_{y}\Big{(}\widehat{p}(x,y,z)-p(x,y,z)\Big{)}|$
$\leq\sum_{xz}\sum_{y}|\widehat{p}(x,w,z)-p(x,w,z)|=\|p_{XYZ}-\widehat{p}_{XYZ}\|\leq\epsilon\,.$
Similarly,
$\|\widehat{p}_{Z}-p_{Z}\|_{1}$ $=\sum_{z}|\widehat{p}(z)-p(z)|=\sum_{z}|\sum_{xy}\Big{(}\widehat{p}(x,y,z)-p(x,y,z)\Big{)}|$
$\leq\sum_{z}\sum_{xy}|\widehat{p}(x,w,z)-p(x,w,z)|=\|p_{XYZ}-\widehat{p}_{XYZ}\|\leq\epsilon\,.$
Since the estimator $\widehat{m}^{B}$ involves the estimation of $p_{XY\,|\,Z}$ and $p_{W\,|\,Z}$ for $W=X$ or $Y$, we proceed to bound the error of them. For any $(x,y,z)\in\{0,1\}^{3}$,
$|\widehat{p}(x,y\,|\,z)-p(x,y\,|\,z)|$ $=\Big{|}\frac{\widehat{p}(x,y,z)}{\widehat{p}(z)}-\frac{p(x,y,z)}{p(z)}\Big{|}$
$=\Big{|}\frac{\widehat{p}(x,y,z)}{\widehat{p}(z)}-\frac{\widehat{p}(x,y,z)}{p(z)}+\frac{\widehat{p}(x,y,z)}{p(z)}-\frac{p(x,y,z)}{p(z)}\Big{|}$
$\leq\widehat{p}(x,y,z)\Big{|}\frac{1}{\widehat{p}(z)}-\frac{1}{p(z)}\Big{|}+\frac{1}{p(z)}|\widehat{p}(x,y,z)-p(x,y,z)|$
$\leq\widehat{p}(x,y,z)\frac{|\widehat{p}(z)-p(z)|}{p(z)\widehat{p}(z)}+\frac{1}{p(z)}|\widehat{p}(x,y,z)-p(x,y,z)|$
$=\frac{1}{p(z)}\Big{(}\widehat{p}(x,y\,|\,z)|\widehat{p}(z)-p(z)|+|\widehat{p}(x,y,z)-p(x,y,z)|\Big{)}$
$\leq\frac{1}{p(z)}\Big{(}|\widehat{p}(z)-p(z)|+|\widehat{p}(x,y,z)-p(x,y,z)|\Big{)}\,.$
Analogously, for $W=X$ or $Y$ and any $(w,z)\in\{0,1\}^{2}$
$|\widehat{p}(w\,|\,z)-p(w\,|\,z)|$ $=\Big{|}\frac{\widehat{p}(w,z)}{\widehat{p}(z)}-\frac{p(w,z)}{p(z)}\Big{|}$
$=\Big{|}\frac{\widehat{p}(w,z)}{\widehat{p}(z)}-\frac{\widehat{p}(w,z)}{p(z)}+\frac{\widehat{p}(w,z)}{p(z)}-\frac{p(w,z)}{p(z)}\Big{|}$
$\leq\widehat{p}(w,z)\Big{|}\frac{1}{\widehat{p}(z)}-\frac{1}{p(z)}\Big{|}+\frac{1}{p(z)}|\widehat{p}(w,z)-p(w,z)|$
$\leq\widehat{p}(w,z)\frac{|\widehat{p}(z)-p(z)|}{p(z)\widehat{p}(z)}+\frac{1}{p(z)}|\widehat{p}(w,z)-p(w,z)|$
$=\frac{1}{p(z)}\Big{(}\widehat{p}(w\,|\,z)|\widehat{p}(z)-p(z)|+|\widehat{p}(w,z)-p(w,z)|\Big{)}$
$\leq\frac{1}{p(z)}\Big{(}|\widehat{p}(z)-p(z)|+|\widehat{p}(w,z)-p(w,z)|\Big{)}\,.$
Denote $\widehat{p}(x,y,z)=p(x,y,z)+\delta_{xyz}$, $\widehat{p}(w,z)=p(w,z)+\delta_{wz}$, $\widehat{p}(z)=p(z)+\delta_{z}$, thus $\delta_{xyz},\delta_{wz}$ are bounded correspondingly. Then we are ready to show the concentration of $\widehat{m}^{B}$ below.
$|\widehat{m}^{B}-m^{B}|$ $=\sum_{z}\bigg{\{}\widehat{p}(z)\sum_{xy}|\widehat{p}(x,y\,|\,z)-\widehat{p}(x\,|\,z)\widehat{p}(y\,|\,z)|-p(z)\sum_{xy}|p(x,y\,|\,z)-p(x\,|\,z)p(y\,|\,z)|\bigg{\}}\,,$
where
$\widehat{p}(z)\sum_{xy}|\widehat{p}(x,y\,|\,z)-\widehat{p}(x\,|\,z)\widehat{p}(y\,|\,z)|$
$=$ $\Big{(}p(z)+\delta_{z}\Big{)}\sum_{xy}\Big{|}\Big{(}p(x,y\,|\,z)+\delta_{xyz}\Big{)}-\Big{(}p(x\,|\,z)+\delta_{xz}\Big{)}\Big{(}p(y\,|\,z)+\delta_{yz}\Big{)}\Big{|}$
$\leq$ $p(z)\sum_{xy}|p(x,y\,|\,z)-p(x\,|\,z)p(y\,|\,z)|+p(z)\sum_{xy}\Big{\{}|\delta_{xyz}|+|\delta_{xz}\delta_{yz}+\delta_{xz}p(y\,|\,z)+\delta_{yz}p(x\,|\,z)|\Big{\}}$
$+\delta_{z}\sum_{xy}|\widehat{p}(x,y\,|\,z)-\widehat{p}(x\,|\,z)\widehat{p}(y\,|\,z)|$
$\leq$ $p(z)\sum_{xy}|p(x,y\,|\,z)-p(x\,|\,z)p(y\,|\,z)|+p(z)\sum_{xy}\Big{\{}|\delta_{xyz}|+|\delta_{xz}\widehat{p}(y\,|\,z)+\delta_{yz}p(x\,|\,z)|\Big{\}}+8\delta_{z}$
$\leq$ $p(z)\sum_{xy}|p(x,y\,|\,z)-p(x\,|\,z)p(y\,|\,z)|+p(z)\sum_{xy}\Big{\{}|\delta_{xyz}|+|\delta_{xz}|+|\delta_{yz}|\Big{\}}+8\delta_{z}\,.$
Therefore,
$|\widehat{m}^{B}-m^{B}|$ $\leq\sum_{z}\bigg{(}p(z)\sum_{xy}\Big{\{}|\delta_{xyz}|+|\delta_{xz}|+|\delta_{yz}|\Big{\}}+8\delta_{z}\bigg{)}$
$\leq 4\|p_{Z}-\widehat{p}_{Z}\|_{1}+\|p_{XYZ}-\widehat{p}_{XYZ}\|_{1}$
$\quad 4\|p_{Z}-\widehat{p}_{Z}\|_{1}+2\|p_{XZ}-\widehat{p}_{XZ}\|_{1}$
$\quad 4\|p_{Z}-\widehat{p}_{Z}\|_{1}+2\|p_{YZ}-\widehat{p}_{YZ}\|_{1}$
$\quad 8\|p_{Z}-\widehat{p}_{Z}\|_{1}$
$=20\|p_{Z}-\widehat{p}_{Z}\|_{1}+\|p_{XYZ}-\widehat{p}_{XYZ}\|_{1}+2\|p_{XZ}-\widehat{p}_{XZ}\|_{1}++2\|p_{YZ}-\widehat{p}_{YZ}\|_{1}$
$\leq 25\epsilon\,.$
Choosing $\epsilon=c/50$, we have with probability at least
$1-3\exp(-C_{0}nc^{2})\,,$
$|\widehat{m}^{B}-m^{B}|\leq c/2$ and $\psi^{B}$ is correct, where the constant $C_{0}=50^{2}\times 25$. ∎
D.2 Proof of lower bound for Bernoulli distribution
Proof.
We proceed to verify the validity of $p_{0}^{B}$ and $p_{1}^{B}$ by showing they satisfy (C2) and (C3). We can compute the KL divergence between them:
$\mathbf{KL}(p_{1}^{B}\|p_{0}^{B})=\log(1-4c^{2})+2c\log(1+\frac{4c}{1-2c})\leq 100c^{2}\,,$
for $c$ small enough. Then (C2) holds. Moreover, it suffices to show for $p_{1}^{B}$ is $c$-strong tree-faithfulness to the chain graph $Z\to X\to Y$. To achieve this, we can compute
$m^{B}(X,Y)=2c,m^{B}(X,Z)=2c,$
$m^{B}(X,Y\,|\,Z)\geq 2c(1-4c^{2})\geq c,$
$m^{B}(Z,X\,|\,Y)\geq 2c(1-4c^{2})\geq c\,,$
for $c$ small enough. Then (C3) holds. ∎
Appendix E Proof of Theorem 4.2 (Gaussian distribution)
Proof.
The upper bound (C1) of $\psi^{G}$ is given by Lemma C.1 in Wang et al. (2024). We proceed to verify the validity of $p_{0}^{G}$ and $p_{1}^{G}$ by showing they satisfy (C2) and (C3). We can compute the KL divergence between them:
$\mathbf{KL}(p_{1}^{G}\|p_{0}^{G})=-\log(1-4c^{2})\leq 100c^{2}\,,$
for $c$ small enough. Then (C2) holds. Moreover, it suffices to show for $p_{1}^{G}$ is $c$-strong tree-faithfulness to the chain graph $Z\to X\to Y$. To achieve this, we can compute
$m^{G}(X,Y)=2c,m^{G}(X,Z)=2c,$
$m^{G}(X,Y\,|\,Z)\geq 2c(1-4c^{2})\geq c,$
$m^{G}(Z,X\,|\,Y)\geq 2c(1-4c^{2})\geq c\,,$
for $c$ small enough. Then (C3) holds. ∎
Appendix F Proof of Theorem 5.1 (nonparametric continuous distribution)
Proof.
The upper bound (C1) of $\psi^{NP}$ is given by Theorem 5.6 in Neykov et al. (2021). We proceed to specify the constructions of $p_{0}^{NP}$ and $p_{1}^{NP}$ and show they satisfy (C2) and (C3).
Let $p_{0}^{NP}$ be independent uniform distributions $Unif^{3}[0,1]$. Thus, $p_{0}^{NP}$ is Markov to an empty graph, and satisfies Lipschitz and smoothness condition. For $p_{1}^{NP}$, we design it to be Markov to a $V$-structure $X\rightarrow Z\leftarrow Y$, which is a three-node poly-forest. Under this graph, a faithful distribution is supposed to have $X\perp\!\!\!\perp Y$ while $X\not\perp\!\!\!\perp Y\,|\,Z$ and $X\not\perp\!\!\!\perp Z,Y\not\perp\!\!\!\perp Z$. In particular, $p_{1}^{NP}$ is specified as follows (we will suppress the superscript and subscript by writing it as $p$ to avoid notation clutter): $X,Y\sim Unif[0,1]$ and $p_{Z\,|\,X,Y}$ is a mixture of perturbation to uniform distribution, whose component depends on independent multi-dimensional Radmacher random variables $\Delta\in\{-1,1\}^{m^{\prime}\times m^{\prime}},\nu\in\{-1,1\}^{m}$ for some positive integers $m,m^{\prime}$ which are specified later. Then
$p(Z\,|\,X,Y)=\mathbb{E}_{\Delta,\nu}\bigg{[}1+$ $\gamma_{\Delta}(X,Y)\eta_{\nu}(Z)\bigg{]}\,,$
where
$\gamma_{\Delta}(x,y)$ $=\rho^{2}\sum_{i\in[m^{\prime}]}\sum_{j\in[m^{\prime}]}\Delta_{ij}h_{ij,m^{\prime}}(x,y)$
$\eta_{\nu}(z)$ $=\rho\sum_{j\in[m]}\nu_{j}h_{j,m}(z)$
$h_{ij,m^{\prime}}(x,y)$ =$=\begin{cases}\sqrt{m^{\prime}}^{2}\widetilde{h}(m^{\prime}x-i+1,m^{\prime}y-j+1)&\forall(x,y)\in[\frac{i-1}{m},\frac{i}{m}]\otimes[\frac{j-1}{m},\frac{j}{m}]\\ 0&\text{otherwise}\end{cases}$
$h_{j,m}(z)$ =$=\begin{cases}\sqrt{m}h(mz-j+1)&\forall z\in[\frac{j-1}{m},\frac{j}{m}]\\ 0&\text{otherwise}\end{cases}$
[…]
for some function $h(x)$ infinitely differentiable on $[0,1]$ such that
[…]
for some $\rho>0$ that will be specified later, and some constants $a,b_{1},b_{2}>0$.
Now we show that each $p$ satisfy the $c$-strong tree-faithfulness and smoothness conditions. Since the operation $\mathbb{E}_{\Delta,\nu}$ is linear, we consider one instance of $(\Delta,\nu)$ in the following discussion.
$c$-strong tree-faithfulness
We highlight several important observations:
$p(z)$ $=\int_{x,y}p(z\,|\,x,y)p(x)p(y)=\int_{x,y}p(z\,|\,x,y)=1$
$p(x\,|\,z)$ $=\frac{p(x,z)}{p(z)}=p(x,z)=\int_{y}p(z\,|\,x,y)=1+\Big{[}\rho^{2}\sum_{i}\Big{(}\sum_{j}\Delta_{ij}\Big{)}h_{i,m^{\prime}}(x)\Big{]}\eta_{\nu}(z)$
$p(y\,|\,z)$ $=\frac{p(y,z)}{p(z)}=p(y,z)=\int_{x}p(z\,|\,x,y)=1+\Big{[}\rho^{2}\sum_{j}\Big{(}\sum_{i}\Delta_{ij}\Big{)}h_{j,m^{\prime}}(y)\Big{]}\eta_{\nu}(z)\,.$
Therefore, tree-faithfulness is satisfied (while strong version still needs to be shown):
$p(x,z)-p(x)p(z)$ $=\Big{[}\rho^{2}\sum_{i}\Big{(}\sum_{j}\Delta_{ij}\Big{)}h_{i,m^{\prime}}(x)\Big{]}\eta_{\nu}(z)\neq 0$
$p(y,z)-p(y)p(z)$ $=\Big{[}\rho^{2}\sum_{j}\Big{(}\sum_{i}\Delta_{ij}\Big{)}h_{j,m^{\prime}}(y)\Big{]}\eta_{\nu}(z)\neq 0$
$p(x,y\,|\,z)-p(x\,|\,z)p(y\,|\,z)$ =$=\bigg{\{}\gamma_{\Delta}(x,y)$
$\quad-\Big{[}\rho^{2}\sum_{j}\Big{(}\sum_{i}\Delta_{ij}\Big{)}h_{j,m^{\prime}}(y)\Big{]}-\Big{[}\rho^{2}\sum_{i}\Big{(}\sum_{j}\Delta_{ij}\Big{)}h_{i,m^{\prime}}(x)\Big{]}\bigg{\}}\eta_{\nu}(z)$
$\quad-\rho^{4}\Big{[}\sum_{i}\Big{(}\sum_{j}\Delta_{ij}\Big{)}h_{i,m^{\prime}}(x)\Big{]}\Big{[}\sum_{j}\Big{(}\sum_{i}\Delta_{ij}\Big{)}h_{j,m^{\prime}}(y)\Big{]}\eta_{\nu}^{2}(z)\neq 0\,.$
By Lemma B.4 in Neykov et al. (2021), it suffices to show the above three nonzero quantities are bounded away from zero in $L_{1}$. Specifically, because we have for any $i$ or $j$,
$0.7\sqrt{m^{\prime}}\leq\mathbb{E}_{\Delta}|\sum_{j}\Delta_{ij}|=\mathbb{E}_{\Delta}|\sum_{i}\Delta_{ij}|\leq\sqrt{m^{\prime}}\,.$
Then
$\|p(x,y\,|\,z)-p(x\,|\,z)p(y\,|\,z)\|_{1}$ $\geq\rho^{2}\int\Big{|}\sum_{i,j}\Delta_{ij}\Big{[}h_{ij,m^{\prime}}(x,y)-h_{i,m^{\prime}}(x)-h_{j,m^{\prime}}(y)\Big{]}\Big{|}\Big{|}\eta_{v}(z)\Big{|}$
$\quad-\rho^{4}\int\Big{|}\Big{[}\sum_{i}\Big{(}\sum_{j}\Delta_{ij}\Big{)}h_{i,m^{\prime}}(x)\Big{]}\Big{[}\sum_{j}\Big{(}\sum_{i}\Delta_{ij}\Big{)}h_{j,m^{\prime}}(y)\Big{]}\Big{|}\Big{|}\eta_{v}^{2}(z)\Big{|}$
$\geq\Big{(}\rho^{2}(m^{\prime})^{2}\times\frac{1}{m^{\prime}}g\Big{)}\times\Big{(}\rho\sqrt{m}a\Big{)}-\rho^{4}\Big{(}m^{\prime}\times\sqrt{m^{\prime}}\times\frac{1}{\sqrt{m^{\prime}}}a\Big{)}^{2}\times\Big{(}\rho^{2}m\Big{)}$
$=\rho^{3}m^{\prime}\sqrt{m}\times ag-\rho^{6}(m^{\prime})^{2}m\times a^{2}\,,$
[…]
$\|p(w,z)-p(w)p(z)\|_{1}$ $=\int\rho^{2}\Big{|}\sum_{i}\Big{(}\sum_{j}\Delta_{ij}\Big{)}h_{i,m^{\prime}}(w)\Big{|}\Big{|}\eta_{\nu}(z)\Big{|}$
$\geq\Big{(}\rho^{2}m^{\prime}\times 0.7\sqrt{m^{\prime}}\times\frac{1}{\sqrt{m^{\prime}}}a\Big{)}\times\Big{(}\rho\sqrt{m}a\Big{)}$
$=\rho^{3}m^{\prime}\sqrt{m}a^{2}\,.$
We will lower bound both of them above by the order of $c$ when specifying the parameters.
Lipschitz condition
We then check the TV distance between $p(x,y\,|\,z)$ and $p(x,y\,|\,z^{\prime})$:
$\|p(x,y\,|\,z)-p(x,y\,|\,z^{\prime})\|_{1}$ $\leq\int\Big{|}\gamma_{\Delta}(x,y)\Big{|}\Big{|}\eta_{\nu}(z)-\eta_{\nu}(z^{\prime})\Big{|}$
$\leq b_{2}m^{\prime}\rho^{2}\times\Big{|}\eta_{\nu}(z)-\eta_{\nu}(z^{\prime})\Big{|}$
$\leq\bigg{[}(b_{2}m^{\prime}\rho^{2})\times(m^{1/2}m\rho\|h^{\prime}\|_{\infty})\bigg{]}|z-z^{\prime}|\,.$
We will need the term in the bracket to be smaller than some constant.
Smoothness condition
We check the Hölder smoothness of $p(x,y\,|\,z)$ in $(x,y)$. Following the proof Theorem 4.2 in Neykov et al. (2021), for any $k\leq\lfloor s\rfloor$
$\bigg{|}\frac{\partial^{k}}{\partial x^{k}}\frac{\partial^{\lfloor s\rfloor-k}}{\partial y^{\lfloor s\rfloor-k}}p(x,y\,|\,z)-\frac{\partial^{k}}{\partial x^{k}}\frac{\partial^{\lfloor s\rfloor-k}}{\partial y^{\lfloor s\rfloor-k}}p(x^{\prime},y^{\prime}\,|\,z)\bigg{|}$
$\leq$ $\bigg{|}\frac{\partial^{k}}{\partial x^{k}}\frac{\partial^{\lfloor s\rfloor-k}}{\partial y^{\lfloor s\rfloor-k}}\gamma_{\Delta}(x,y)\eta_{\nu}(z)-\frac{\partial^{k}}{\partial x^{k}}\frac{\partial^{\lfloor s\rfloor-k}}{\partial y^{\lfloor s\rfloor-k}}\gamma_{\Delta}(x^{\prime},y^{\prime})\eta_{\nu}(z)\bigg{|}$
$\leq$ $m^{1/2}\rho\|h\|_{\infty}\bigg{|}\frac{\partial^{k}}{\partial x^{k}}\frac{\partial^{\lfloor s\rfloor-k}}{\partial y^{\lfloor s\rfloor-k}}\gamma_{\Delta}(x,y)-\frac{\partial^{k}}{\partial x^{k}}\frac{\partial^{\lfloor s\rfloor-k}}{\partial y^{\lfloor s\rfloor-k}}\gamma_{\Delta}(x^{\prime},y^{\prime})\bigg{|}$
$\leq$ $(\rho m^{1/2}a)\times\Big{(}\rho^{2}(m^{\prime})^{s}\sqrt{m^{\prime}}^{2}\bigg{\|}\frac{\partial^{k}}{\partial x^{k}}\frac{\partial^{\lfloor s\rfloor-k}}{\partial y^{\lfloor s\rfloor-k}}\widetilde{h}(x,y)\bigg{\|}_{\infty}\Big{)}\,.$
Thus, the quantity that need to be bounded above by constant is
$\rho^{3}m^{1/2}{m^{\prime}}^{1+s}\,.$
Parameter choice
All in all, we need the choice of $m,m^{\prime},\rho$ to satisfy the following requirements:
  • $c$-strong tree-faithfulness:
    $\rho^{3}m^{1/2}m^{\prime}-\rho^{6}m(m^{\prime})^{2}$ $\gtrsim c$
    $\rho^{3}m^{1/2}{m^{\prime}}$ $\gtrsim c$
  • Lipschitz condition:
    $\rho^{3}m^{3/2}m^{\prime}$ $\lesssim 1$
  • Smoothness condition:
    $\rho^{3}m^{1/2}{m^{\prime}}^{1+s}$ $\lesssim 1$
By setting $m^{\prime}=m^{1/s},\rho^{3}\asymp m^{-(3/2+1/s)},m\asymp c^{-1}$ and assuming $c$ is sufficiently small, above requirements are satisfied. Thus, $p$ falls into the considered model class.
KL divergence
We start by bounding the $\chi^{2}$ divergence then upper bound KL divergence by $\chi^{2}$ divergence:
$\chi^{2}(p_{1}^{NP}\|p_{0}^{NP})+1=$ $\int\frac{(p_{1}^{NP})^{2}}{p_{0}^{NP}}=\mathbb{E}_{\Delta,\Delta^{\prime},\nu,\nu^{\prime}}\int\bigg{[}1+\gamma_{\Delta}(x,y)\eta_{\nu}(z)\bigg{]}\times\bigg{[}1+\gamma_{\Delta^{\prime}}(x,y)\eta_{\nu^{\prime}}(z)\bigg{]}\,.$
The integral is
$\int\bigg{[}1+\gamma_{\Delta}(x,y)\gamma_{\Delta^{\prime}}(x,y)\eta_{\nu}(z)\eta_{\nu^{\prime}}(z)\bigg{]}=1+\rho^{6}\langle\Delta,\Delta^{\prime}\rangle\langle\nu,\nu^{\prime}\rangle\,.$
Therefore,
$\chi^{2}(p_{1}^{NP}\|p_{0}^{NP})+1$ $=\mathbb{E}_{\Delta,\Delta^{\prime},\nu,\nu^{\prime}}\bigg{\{}1+\rho^{6}\langle\Delta,\Delta^{\prime}\rangle\langle\nu,\nu^{\prime}\rangle\bigg{\}}\leq\mathbb{E}_{\Delta,\Delta^{\prime},\nu,\nu^{\prime}}\bigg{\{}\exp\bigg{(}\rho^{6}\langle\Delta,\Delta^{\prime}\rangle\langle\nu,\nu^{\prime}\rangle\bigg{)}\bigg{\}}\,.$
Following the proof Theorem 4.2 in Neykov et al. (2021), we can upper bound the right hand side above and obtain
$\chi^{2}(p_{1}^{NP}\|p_{0}^{NP})+1\leq\sqrt{\frac{1}{1-(\rho^{6})^{2}m{m^{\prime}}^{2}}}\,.$
Since the function $f(t)=1/\sqrt{1-t^{2}}\leq 2t+1$ for $t$ sufficiently small, with the choice of $\rho,m,m^{\prime}$, we have for sufficiently small $c$,
$\chi^{2}(p_{1}^{NP}\|p_{0}^{NP})+1\leq C_{0}\times c^{\frac{5s+2}{2s}}+1\,,$
for some constant $C_{0}$. Therefore, we arrive at
$\mathbf{KL}(p_{1}^{NP}\|p_{0}^{NP})\leq\chi^{2}(p_{1}^{NP}\|p_{0}^{NP})\lesssim c^{\frac{5s+2}{2s}}$
Application of Corollary G.4 completes the proof. ∎
Appendix G Auxiliary lemmas
For lower bound techniques, we mainly apply the Fano’s inequality and Tsybokov’s method.
Lemma G.1 (Yu (1997), Lemma 3).
For a model family $\mathcal{M}$ contains $M$ many distributions indexed by $j=1,2,\ldots,M$ such that
$\alpha$ $=\max_{P_{j}\neq P_{k}\in\mathcal{M}}\mathbf{KL}(P_{j}\|P_{k})$
$s$ $=\min_{P_{j}\neq P_{k}\in\mathcal{M}}\mathbf{dist}(\theta(P_{j}),\theta(P_{k}))\,,$
where $\theta$ is a functional of its distribution argument. Then for any estimator $\widehat{\theta}$ for $\theta(P)$,
$\inf_{\widehat{\theta}}\sup_{P\in\mathcal{M}}\mathbb{E}_{P}\mathbf{dist}(\theta(P),\widehat{\theta})\geq\frac{s}{2}\bigg{(}1-\frac{\alpha+\log 2}{\log M}\bigg{)}\,.$
Lemma G.2 (Tsybakov (2008), Theorem 2.5).
For a model family $\mathcal{M}$ contains distributions $P_{0},P_{1},\ldots,P_{M}$ with $M\geq 2$ and suppose that $\Theta$ contains elements $\theta_{0},\theta_{1},\ldots,\theta_{M}$ such that:
  1. 1.
    $\mathbf{dist}(\theta_{j},\theta_{k})\geq 2s$, $\forall 0\leq j<k\leq M$;
  2. 2.
    $P_{j}\ll P_{0}$, $\forall j=1,\ldots,M$, and
    $\frac{1}{M}\sum_{j=1}^{M}\mathbf{KL}(P_{j},P_{0})\leq\alpha\log M$
    with $80<\alpha<1/$ 8 and $P_{j}=P_{\theta_{j}}$. Then
$\inf_{\widehat{\theta}}\sup_{\theta\in\Theta}\mathbb{P}_{\theta}\Big{(}\mathbf{dist}(\theta,\widehat{\theta})\geq s\Big{)}\geq\frac{\sqrt{M}}{1+\sqrt{M}}\bigg{(}1-2\alpha-\frac{2\alpha}{\log M}\bigg{)}\,.$
Set $\theta(P_{j})=j$ to be the index, ${}\mathbf{dist}(\cdot,\cdot)=\mathbf{1}\{\cdot\neq\cdot\}$, consider $P_{j}$ to be a product measure of $n$ i.i.d. samples for any $P_{j}\in\mathcal{M}$, then Lemma G.1 and G.2 under model selection context can be stated as follows:
Corollary G.3 (Fano’s inequality).
For a model family $\mathcal{M}$ contains $M$ many distributions indexed by $j=1,2,\ldots,M$ such that $\alpha=\max_{P_{j}\neq P_{k}\in\mathcal{M}}\mathbf{KL}(P_{j}\|P_{k})$. If the sample size is bounded as
$n\leq\frac{(1-2\delta)\log M}{\alpha}\,,$
then for any estimator $\widehat{\theta}$ for the model index:
$\inf_{\widehat{\theta}}\sup_{j\in[M]}P_{j}(\widehat{\theta}\neq j)\geq\delta-\frac{\log 2}{\log M}\,.$
Corollary G.4 (Tsybakov’s method).
For a model family $\mathcal{M}$ contains $M$ many distributions indexed by $j=1,2,\ldots,M$ such that $\alpha=\max_{j\in[M]}\mathbf{KL}(P_{j}\|P_{0})$. If the sample size is bounded as
$n\leq\frac{\log M}{16\alpha}\,,$
then for any estimator $\widehat{\theta}$ for the model index:
$\inf_{\widehat{\theta}}\sup_{j\in[M]}P_{j}(\widehat{\theta}\neq j)\geq\frac{1}{16}\,.$
Appendix H Experiment details
We describe the experiment details of Section 6 and provide additional results in this appendix. The code to replicate the experiments is available at https://anonymous.4open.science/r/OptimalPolyforests-5F64/README.md.
Graph generation
For our experiments, we simulate poly-forests by initializing an empty adjacency matrix, then randomly ordering the nodes. For each node along the ordering (except the first), an edge is added from a random preceding node in the ordering with probability 80%, ensuring acyclicity and forming a directed forest.
Gaussian distribution
We simulate random Gaussian poly-forests according to the following structural equation model:
$X_{k}=\beta_{k}\times X_{\operatorname{pa}(k)}+\eta_{k}\,,\quad\eta_{k}\sim\mathcal{N}(0,\sigma_{k}^{2}),$
where the coefficients are sampled as $\beta_{k}\sim Unif([-0.5,-0.1]\cup[0.1,0.5])$, and the noise variances are fixed at $\sigma^{2}_{k}\equiv 1$. Under the Gaussian assumption, we test conditional independence using partial correlations (4). We set the cutoff to be 0.05.
Bernoulli distribution
For the synthetic Bernoulli data, root nodes are sampled independently from a $Bern(0.5)$. For each non-root node $X_{k}$, its conditional distribution given its parents $X_{\operatorname{pa}(k)}$ is given as follows. Let $b_{k}\sim Unif(l,u)$ and $R_{k}\sim Unif\{-1,1\}$, and define the conditional probabilities by:
$X_{k}\,|\,X_{\operatorname{pa}(k)}=1\sim Bern(0.5+R_{k}\times b_{k})$
$X_{k}\,|\,X_{\operatorname{pa}(k)}=0\sim Bern(0.5-R_{k}\times b_{k})\,.$
This construction introduces parent-dependent dependence while ensuring the conditional probabilities remain within a valid range. We set $l=0.3,u=0.48$ in our experiments. In this Bernoulli case, we employ the test (4.1) with the cutoff being 0.05.
Nonparametric continuous distribution
To generate synthetic nonparametric continuous data, root nodes are drawn from Uniform distribution $U(0,1)$. Each subsequent node $X_{k}$ is generated as a weighted sum of nonlinear transformation of the parent and an individual uniform noise:
$X_{k}=f_{k}(X_{\operatorname{pa}(k})\times\frac{3}{10}+U_{k}(0,1)\times\frac{7}{10}$
where $U_{k}(0,1)\sim Unif(0,1)$. The transformation functions $f_{k}(z)$ are chosen uniformly at random for each parent-child link from a predefined set below, including functions with both range and domain being $[0,1]$. This process introduces nonparametric dependencies.
$f_{k}(z)\sim Unif\bigg{\{}0.5\times\Big{[}\sin(2\pi z)+1\Big{]},z^{2},\frac{\log(1+z)}{\log 2},0.5\times\Big{[}\cos(2\pi z)+1\Big{]}\bigg{\}}$
For continuous data, we first apply a discretization procedure to convert real-valued observations into categorical representations suitable for contingency-table-based analysis. Following the suggestion of Theorem 5.6 in Neykov et al. (2021), we partition the range into a number of bins determined by the smoothness parameter and the sample size. For variables $(X,Y)$, we use $n^{2/(5s+2)}$ number of bins. For variable $Z$, we use $n^{2s/(5s+2)}$ number of bins. We set $s=1$ in this experiment. We assign each observation to a discrete bin in a two- or three-way contingency table, respectively.
Once discretized, we apply the U-statistic Conditional Independence (UCI) test (Kim et al., 2024) to assess whether $X\perp\!\!\!\perp Y\,|\,Z$. The test computes a U-statistic within each stratum defined by a unique bin of the conditioning variable $Z$, and aggregates the results across strata. We apply weighted U-statistic in our empirical experiments. Since UCI is a permutation test, we set the number of permutation to be 199, following the code provided in https://github.com/ilmunk/UCI. We again set the cutoff to be 0.05.
Figure 2: Precise Recovery Rate (PRR) vs. sample size for poly-forest learning for Bernoulli, Gaussian, and nonparametric continuous distributions over varying number of nodes (indicated by colors). PRR consistently increase toward 100% as sample size increases across all experimental settings.
Evaluation
For each experiment setup, we report the average (over 50 random instantiations) Structural Hamming Distance (SHD) between the ground truth and our estimated graph skeleton in Figure 1, we also evaluate the Precise Recovery Rate (PRR), with the results reported in Figure 2. PRR measures the relative frequency of exact recovery of the true graph structure. Our experimental results demonstrate the robust performance of the PC-tree algorithm across various data distributions. As illustrated in the provided subplots for Gaussian, Bernoulli, and Nonparametric synthetic data, the PRR consistently converges towards 100% with increasing sample size. Our empirical results support the theoretical guarantees of the PC-tree algorithm.
Experiment setting
We consider number of nodes $d=[20,40,60,80,100]$. To demonstrate the convergence of SHD toward zero, we vary the sample size $n$ from 300 to 3000 across all data types (Bernoulli, Gaussian, and nonparametric) in Figure 1, enabling a consistent evaluation of structural accuracy as sample size increases. To examine convergence behavior in PRR plots, we vary the sample size according to the data type: for nonparametric data, sample sizes range from 700 to 2000; for Bernoulli data, from 3000 to 7000; and for Gaussian data, from 1000 to 10000. The difference is for better presentation and comes from the signal contained in each distribution setup.
Compute resources
All experiments were conduced on an Intel Core i7-12800H 2.40GHz CPU.